Determine a)the capacitance and bthe maximum potencial difference that can be applied to a teflon-filled parallel-plate capacitor having a olate area of 1.75 cm^2 and a plate seperation of 0.040 0 mm.
1
Expert's answer
2020-04-02T10:42:01-0400

a)


"C=
ε_
r
ε_
0\frac{A}{d}"

"C=2.1(8.854\cdot10^{-12})\frac{1.75\cdot10^{-4}}{4\cdot10^{-5}}=8.13\cdot10^{-11}F"

b)


"U=Ed=(60\cdot10^{6})(4\cdot10^{-5})=2400\ V=2.4\ kV"